mathslover (mathslover):

prove s = ut + 1/2 at^2

mathslover (mathslover):

Prove that s = ut + \(\frac{1}{2}at^2\) : v = \(\frac{s}{t}\) u +at = \(\frac{s}{t}\) ut + \(at^2\) = s

mathslover (mathslover):

? where did i do wrong?

mathslover (mathslover):